PROGRESS REPORT ON A PORTABLE TI:SAPPHIRE COMB LASER WITH FREQUENCIES REFER-RING TO CESIUM ATOM TWO-PHOTON TRANSITIONS

\underline{WANG-YAU CHENG}, CHIEN-MING WU, TZ-WEI LIU, YO-HUAN CHEN, IAMS, Academia Sinica, No. 1, Roosevelt Rd., Sec. 4, Taipei, Taiwan.

A portable Ti:sapphire comb laser would contribute significantly to generalize comb-laser applications, such as the astrocomb missions or other interdisciplinary collaborations. To develop a portable comb laser, three barriers lie ahead: one is to miniaturize and robotize the frequency reference system of the comb laser; the second is to ensure the long-term frequency accuracy without satellite connection, and the third is to miniaturize the pumping laser system. We developed two hand-size cesium-stabilized diode lasers at 822 nm and 884 nm to serve as frequency references for a comb laser and we carried out a comb-laser-based CPT experiment with one single cesium cell that might offer a locking procedure for long-term comb laser accuracy. We will also report our plans and progress on a fiber laser pumped Ti:sapphire comb laser.